# A Two-Stage Stochastic Model for Road-Rail Intermodal Freight Transportation Under Demand and Capacity Uncertainty

### v1

@@ 3 Problem Description
We address the optimization challenge faced by transportation and manufacturing companies managing intermodal logistics involving predefined and fixed train schedules. These entities operate multiple warehouses and/or manufacturing plants, with the goal of transporting containers from these facilities to intermodal or train stations in time for train departures. The utilization of spot capacities available on trains at intermodal hubs is particularly crucial in such operations. These capacities are often subject to significant randomness due to fluctuating demand and competing priorities from other users, such as long-term contracts and operational constraints (Wang and Meng,, 2021). Addressing these challenges requires robust planning and optimization strategies tailored to intermodal logistics.
The primary focus of this study involves intermodal transportation using two modes: rail (trains) and road (trucks). Containers must be transported by road from various supply locations to specific intermodal hubs, ensuring they meet the fixed schedules of departing trains. Each train operates on a predetermined route and schedule, with the company utilizing the available capacity to transport containers to their respective destinations. The demand for container transport is defined for each train, aligning with the company’s logistics requirements. Given that only one train typically serves a particular destination within a specified period, the optimization process focuses on allocating resources efficiently to meet these constraints.
A distinguishing feature of this problem is the inherent stochasticity of both demand and train capacities. Demand variability reflects real-world logistics scenarios, where requirements at destination nodes are often uncertain (Demir et al.,, 2016). Additionally, the exact train capacities available at intermodal hubs are unpredictable (Lu et al.,, 2020), as they are influenced by shared usage among multiple users, including long-term contract commitments and operational constraints. Similar challenges are observed in container shipping, where fluctuating demand and capacity uncertainties significantly impact operational decisions (Wang and Meng,, 2021; Nurminarsih et al.,, 2024). Addressing these uncertainties necessitates adaptive planning approaches capable of responding dynamically to real-time variations, ensuring both logistics reliability and cost-effectiveness.
The problem requires companies to first determine the number of containers to prepare at supply locations, such as warehouses or manufacturing plants, with an associated preparation cost. Containers are considered ready for shipment as-is, regardless of whether they are fully or partially loaded, once deemed ready for order fulfillment. The next step involves determining the optimal transportation plan, including the selection of supply locations and intermodal hubs, to minimize costs and meet demand at destination nodes within the specified time windows.
To make this discussion more relatable, consider a practical example from our case study. A freight train route from Seattle, WA, to Chicago, IL, involves intermediate stops at Spokane, WA; Billings, MT; Fargo, ND; and Minneapolis, MN. The company must decide which containers from nearby facilities should be sent to each intermediate hub to align with the train’s fixed schedule. Given limited capacity at each hub due to competing cargo, this decision involves balancing demand at the destination with the availability of containers at supply locations, all while accounting for potential delays and environmental considerations.
Key constraints include the fixed train schedules, stochastic capacity availability at each station, and the specified delivery time windows for shipments. Containers must be transported via road to align with train schedules, optimizing the assignment of shipments to trains while minimizing penalties incurred from unmet demand. Environmental considerations are also integral to the problem, with the objective of reducing carbon emissions by selecting the closest supply locations for shipments, thereby minimizing overall travel distances.
Routing decisions from the supply/origin locations to the intermodal hubs are excluded from this problem. This exclusion assumes that these trips are carried out frequently using pre-determined, established routes. For companies managing intermodal logistics, such routes are often fixed due to operational consistency and infrastructure constraints. Treating these routes as fixed simplifies the problem while remaining realistic for practical applications. However, incorporating dynamic routing considerations could be a valuable extension in scenarios where flexibility is necessary, such as changing traffic conditions or infrastructure disruptions. We leave this as a direction for future research.
By integrating these considerations, the problem focuses on designing a transportation plan that minimizes costs, aligns with operational constraints, and reduces environmental impact. The utilization of spot capacities at intermodal hubs ensures efficient resource allocation despite the inherent uncertainties in train schedules and capacities. This adaptive approach addresses real-world logistical complexities, making it applicable to diverse operational settings. We consider the following assumptions in our model:
  • Containers are shipped as-is once deemed ready for order fulfillment, regardless of whether they are fully or partially loaded, ensuring operational simplicity.
  • Supply locations are pre-loaded with containers ready for transport, and delays or costs associated with procuring empty containers are not considered.
  • Train routes, schedules, and travel times are fixed and known during the planning horizon. Manufacturing and logistics companies have no control over these decisions, as they are determined by rail operators and infrastructure constraints.
  • Traffic conditions and potential delays are incorporated into travel time estimates to enhance the robustness of the transportation plan.
  • Containers are uniform in size and weight, simplifying loading and transportation logistics.
  • Once loaded onto a train, containers remain onboard until they reach the designated demand point.
  • Routing decisions from supply locations to intermodal hubs are excluded, as these routes are assumed to follow established patterns typical of many-to-many distribution networks in intermodal transport, where OD pairs often utilize consistent and direct transport paths (Basallo-Triana et al.,, 2021).
This problem description highlights the complex interplay between operational constraints, environmental considerations, and uncertainty in intermodal transportation. Our structured optimization framework ensures adaptability to real-world logistics challenges by integrating spot capacities and demand uncertainty. The next section introduces the mathematical foundation of the model.
@@ 4 Optimization Model
This section presents a two-stage stochastic optimization model designed to address intermodal transportation challenges, particularly uncertainties in demand and train capacities. The model integrates a Conditional Value-at-Risk (CVaR) framework to balance cost efficiency and risk management. Key model components, including sets, parameters, and variables, are outlined in Table 1. The formulation is explained alongside a discussion of constraints, and a deterministic equivalent of the stochastic model is provided.
@@
Set Definition
$I$ Origin locations (warehouses/manufacturing centers)
$J$ Train stations/intermodal hubs
$T$ Time periods
$\Omega$ Scenarios
$N$ Trains
@@
Variable Definition
$x_{ijn\omega t}$ Number of containers shipped from origin $i$ to station $j$ for train $n$ at time $t$ in scenario $\omega$
$y_{i}$ Number of containers prepared at origin $i$
$U_{n\omega}$ Unmet demand for train $n$ in scenario $\omega=\max(0,D_{n\omega}-\mathcal{I}_{j_{f}n\omega})$
@@
$z_{ijnst}$ Binary variable indicating if a container is shipped from origin $i$ to station $j$ for train $n$ at time $t$ in scenario $\omega$
The two-stage structure is well-suited for handling stochastic factors in intermodal transportation. It enables informed initial decisions while accounting for second-stage adjustments based on realized demand and capacity constraints. This approach optimizes container preparation at supply nodes while deferring detailed allocation and fulfillment decisions until more information becomes available, improving adaptability in uncertain conditions.
Stochastic optimization has been widely applied in intermodal logistics for managing demand and capacity uncertainty. Previous studies Tolooie et al., (2020); Huang, (2021) have demonstrated the effectiveness of two-stage models in minimizing costs while responding to real-time disruptions. By leveraging this framework, our model ensures cost-efficient and resilient decision-making in dynamic logistics environments.
In our proposed model, we introduce a two-stage stochastic programming framework that integrates risk aversion. In this framework, both the objective function and the constraints in the second stage incorporate stochastic parameters in line with acceptable practice. Specifically, we adopt a mean-risk objective function that incorporates CVaR with $\alpha\in(0,1)$ being the confidence level. This choice of risk measure allows us to model the risk preferences of decision makers concerning the uncertain total cost, offering flexibility to accommodate various risk thresholds from no risk ($\alpha=0$) to worst-case scenarios ($\alpha\rightarrow 1$).
@@
The objective function in (1) aims to minimize total transportation cost from origins to train stations while also penalizing unmet demand for each train in every scenario. This dual-purpose objective ensures efficient logistics operations by optimizing transportation routes and mitigating the impact of any shortfalls in meeting the demand for each train in every scenario.
Choices are typically made before uncertainties are realized, with recourse decisions deferred to the second stage that addresses the stochasticity. Once the uncertainty is known, detailed second-stage distribution decisions can be addressed. Thus, first-stage logistics planning concentrates on the number of containers prepared at each facility, while the second-stage problem manages demand and capacity uncertainties. We now present the formulation for our two-stage stochastic road-rail intermodal transportation model.
@@ 4.1 Conditional Value-at-Risk (CVaR)
Before delving into the full stochastic optimization model, we provide an overview of the CVaR risk measure, which plays a critical role in its formulation. Initially introduced by Rockafellar et al., (2000) for financial portfolio optimization, CVaR has gained widespread adoption across various fields due to its ability to quantify and mitigate risks associated with extreme outcomes. For example, Künzi-Bay and Mayer, (2006) reformulated CVaR minimization problems as two-stage recourse models in stochastic programming, while Fábián, (2008) developed polyhedral representations to efficiently handle CVaR objectives and constraints in large-scale problems. For further insights into these foundational studies, see Rockafellar et al., (2000); Künzi-Bay and Mayer, (2006); Fábián, (2008).
CVaR measures the expected value of losses exceeding a specified Value-at-Risk (VaR) threshold, making it particularly suitable for capturing "tail risk" in uncertain scenarios. Based on the work of Rockafellar et al., (2000), CVaR can be mathematically expressed as:
$\text{CVaR}_{\alpha}(R)=\min\left\{\text{VaR}+\frac{1}{1-\alpha}\mathbb{E}([R-\text{VaR}]_{+})\mid\text{VaR}\in\mathbb{R}\right\}$ (2)
This holds for $R$ been a random variable, $\alpha\in(0,1)$ denotes the confidence level of the CVaR associated with the expected value of $R$ given that it surpasses the value of $\alpha$-quantile which is the threshold value of (2). The VaR threshold represents the $\alpha$-quantile of $R$ such that the probability of the loss exceeding this value is $1-\alpha$. The notation $[m]_{+}$ = $m\in\mathbb{R}_{+}$ captures non-negative values. To enhance computational efficiency, (2) can be reformulated as a linear program based on the work of Rockafellar et al., (2000) and further explained by Elçi and Noyan, (2018). We adopt this equivalent linear program formulation which is:
@@
$\text{CVaR}_{\alpha}(R)=\min\left\{\theta+\frac{1}{1-\alpha}\sum_{\omega\in\Omega}p_{\omega}\pi_{\omega}\mid\pi_{\omega}\geq y_{\omega}-\theta,\;\forall\,\omega\in\Omega,\;\pi\in\mathbb{R}^{n}_{+},\;\theta\in\mathbb{R}_{+}\right\}$ (3)
CVaR is recognized as a coherent risk measure, satisfying fundamental principles such as monotonicity, subadditivity, and translation equivariance Artzner et al., (1999). It is also classified as a spectral risk measure (Acerbi,, 2002), providing an interpretation as a weighted average of adverse outcomes. The dual representation of CVaR highlights this characteristic:
[…]
This dual form presented in (4) illustrates the ability of CVaR to focus on the most extreme outcomes within the distribution of $R$, offering a robust framework for balancing risk and operational efficiency. Its practical relevance in intermodal transportation lies in its capacity to mitigate disruptions caused by demand variability and capacity uncertainties, as explored in the subsequent sections.
@@ 4.2.1 Objective
Min $\left(\sum_{i\in I}O_{i}y_{i}+(1-\lambda)\mathbb{E}\left(S^{\prime\prime}(a,b,\zeta(\omega),\gamma(\omega)))\right)+\lambda\text{CVaR}_{\alpha}\left(S^{\prime\prime}(a,b,\zeta(\omega),\gamma(\omega))\right)+\sum_{t\in T}\rho\eta_{\omega}\right)$ (5)
The objective function (5) integrates both the preparation (first-stage) and operational costs (second-stage) for the intermodal stochastic optimization model. This objective minimizes the mean-risk function related to the total stochastic cost, which includes:
  • The operational cost of preparing containers at origins, represented by ($\sum_{i\in I}O_{i}y_{i}$)
  • The expected transportation cost from origins to train stations, and penalties for unmet demand across all scenarios, $(1-\lambda)\mathbb{E}\left(S^{\prime\prime}(a,b,\zeta(\omega),\gamma(\omega))\right)$,
  • The CVaR at confidence level $\alpha$ of the random total cost, $\lambda\text{CVaR}_{\alpha}\left(S^{\prime\prime}(a,b,\zeta(\omega),\gamma(\omega)))\right)$ which accounts for extreme cost realizations,
  • Penalties for exceeding emissions limits, captured by $\sum_{t\in T}\rho\eta_{\omega}$.
This comprehensive objective ensures efficient logistics operations by balancing cost minimization, risk aversion, and sustainability goals, enabling adaptive decision-making under uncertain demand and train capacity.
@@ 4.2.2 Constraints
The constraints applied to the model effectively determine demand assignments, the number of containers on each train per station, and limit the carbon emissions during the planning period. To start with, constraints (6) ensure that the total number of containers transported from each origin $i$ does not exceed the available supply at the origin. Constraints (7) ensure that the number of containers transported to each train station does not exceed the available capacity on each train at a particular station. Further, constraints (8) ensure that container flows $x_{ijn\omega t}$ are linked to binary decision variables $z_{ijn\omega t}$. These constraints enforce that containers can only be shipped if the corresponding binary variable is activated ($z_{ijn\omega t}=1$). The large constant $M$ acts as an upper bound, preventing any flow $x_{ijn\omega t}$ unless the route is selected.

### v2

@@ 3 Problem Description
We address the optimization challenges faced by transportation and manufacturing firms managing intermodal logistics under fixed train schedules. These firms must transport containers from multiple supply points such as warehouses or plants to intermodal hubs in time for scheduled train departures. A critical element of this process is the use of spot capacities on trains, which are often uncertain due to fluctuating demand and competing allocations (e.g., long-term contracts, operational constraints) (Wang and Meng,, 2021).
This study focuses on a two-mode intermodal system (road and rail) where containers are first transported by truck to intermodal hubs, then transferred to trains operating on fixed routes and schedules. Each train serves specific destinations, and the company must align container preparation and dispatch with train departures. The optimization problem involves determining how many containers to prepare at each supply location and how to allocate them across hubs, aiming to meet destination demand on time and at minimal cost.
A core complexity lies in the stochastic nature of both demand and train capacity. Demand at destination nodes vary, and the space on each train is uncertain, often influenced by external users and unpredictable operational factors (Demir et al.,, 2016; Lu et al.,, 2020; Wang and Meng,, 2021). These uncertainties demand adaptive strategies that can dynamically respond to changing conditions while ensuring logistical efficiency and cost-effectiveness.
To illustrate this complexity, consider a real-world example. A freight train travels from Seattle, WA, to Chicago, IL, with scheduled stops at Spokane, WA; Billings, MT; Fargo, ND; and Minneapolis, MN. Multiple supply facilities located near these hubs must decide how many containers to prepare and where to route them. Unused capacity leads to inefficiencies, while over-preparation may create bottlenecks. Compounding this is the need to synchronize road transport with fixed train departures, accounting for varying travel times and uncertain spot availability. These dynamics make planning a multi-layered, high-stakes decision process.
Key constraints include: fixed train schedules, uncertain capacity at hubs, and time-sensitive delivery windows. Companies must prepare containers (regardless of load completeness) and transport them via predetermined road routes to the hubs. While routing decisions between supply points and hubs are excluded from this model, companies consider such routes are often fixed due to operational consistency based on their frequent, standardized nature in practice (Basallo-Triana et al.,, 2021). Environmental impact is another important dimension. The model seeks to reduce emissions by prioritizing nearby supply locations, minimizing road distance and aligning with policy frameworks such as the Clean Air Act (Agency,, 2024) and state-level carbon reduction targets (Institute,, 2024).
Overall, the problem is defined by the need to balance cost, service reliability, and sustainability under high uncertainty. We adopt the following assumptions, grounded in operational norms and common practice: i) Containers are shipped as-is once ready, regardless of load completeness, reflecting industry practices where dispatch cannot be delayed for full consolidation; ii) supply locations are assumed to be pre-stocked with containers, isolating logistics planning from inventory sourcing; iii) train routes and schedules, are treated as fixed, predefined by carriers and not subject to shipper control; iv) road travel times between supply points and hubs considers traffic conditions, suitable for strategic planning where high-frequency variations are averaged out; v) containers are uniform in size, per ISO standards, but may vary in weight due to partial loading vi) containers remain onboard until their final destination, avoiding mid-route handling, reflecting common linehaul practices; and vii) road routing between supply points and hubs is predetermined, consistent with fixed or established distribution protocols (Basallo-Triana et al.,, 2021)
This problem setup reflects the real-world complexity of intermodal freight systems, shaped by uncertain capacities, tight scheduling, and environmental constraints. The following section presents a two-stage stochastic optimization model that captures these dynamics and supports adaptive, cost-effective planning.
@@ 4 Optimization Model
This section presents a two-stage stochastic optimization model designed to address intermodal transportation challenges, particularly uncertainties in demand and train capacities. The model integrates a CVaR framework to balance cost efficiency and risk management. Key model components, including sets, parameters, and variables, are outlined in Table 1. The formulation is explained alongside a discussion of constraints, and a deterministic equivalent of the stochastic model is provided.
@@
Sets and Indices Definition
II $(i)$ Origin locations (warehouses/manufacturing centers)
$J$ $(j)$ Train stations/intermodal hubs
$T$ $(t)$ Time periods
$\Omega$ $(\omega)$ Scenarios
NN $(n)$ Trains
@@ II $(i)$ Origin locations (warehouses/manufacturing centers)
Decision Variable Definition
$x_{ijn\omega t}$ Containers shipped from origin $i$ to station $j$ for train $n$ at time $t$ in scenario $\omega$
$y_{i}$ Containers prepared at origin $i$
$z_{ijn\omega t}$ Binary variable; equals 1 if a container is assigned from $i$ to $j$ for train $n$ at time $t$ in scenario $\omega$, 0 otherwise
Auxiliary Variable Definition
$U_{n\omega}$ Unmet demand for train $n$ in scenario $\omega$
@@
In our proposed model, we introduce a two-stage stochastic programming framework that integrates risk aversion, where both the objective function and second stage constraints incorporate stochastic parameters in line with acceptable practice. Specifically, we adopt a mean-risk objective that incorporates CVaR with $\alpha\in(0,1)$ being the confidence level. This choice of risk measure allows us to model the risk preferences of decision makers concerning the stochastic cost, offering flexibility to accommodate various thresholds from no risk ($\alpha=0$) to worst-case scenarios ($\alpha\rightarrow 1$).
@@
The objective function in (1) minimizes total transportation cost from origins to train stations and penalized unmet demand for each train in every scenario. This dual-purpose objective ensures efficient logistics operations by optimizing transportation routes and mitigating the impact of any shortfalls in meeting the demand for each train in every scenario.
Choices are typically made before uncertainties are realized, with recourse decisions deferred to the second stage that addresses the stochasticity. Once the uncertainty is known, detailed second-stage distribution decisions can be addressed. Thus, first-stage logistics planning concentrates on the number of containers prepared at each facility, while the second-stage problem manages demand and capacity uncertainties.
@@ 4.1 Conditional Value-at-Risk (CVaR)
Before presenting the full stochastic optimization model, we briefly review the CVaR risk measure, which plays a central role in our formulation. Originally introduced by Rockafellar et al., (2000) for portfolio optimization, CVaR has been widely adopted in operations research and logistics due to its ability to quantify the risk of extreme outcomes beyond the traditional Value-at-Risk (VaR) threshold. Applications of CVaR in stochastic programming include two-stage recourse formulations (Künzi-Bay and Mayer,, 2006) and efficient polyhedral approximations for large-scale optimization problems (Fábián,, 2008). For better understanding see (Künzi-Bay and Mayer,, 2006; Fábián,, 2008)
At a given confidence level $\alpha\in(0,1)$, CVaR measures the expected value of losses exceeding the corresponding VaR threshold. Mathematically, CVaR can be defined as:
$\text{CVaR}_{\alpha}(R)=\min_{\text{VaR}\in\mathbb{R}}\left\{\text{VaR}+\frac{1}{1-\alpha}\mathbb{E}\left([R-\text{VaR}]_{+}\right)\right\}$ (2)
Here, $R$ is a random variable representing the loss, and $m]_{+}=\max(m,0)$ denotes the positive part of $m$. The VaR represents the $\alpha$-quantile of $R$, and CVaR captures the average loss beyond this quantile. Intuitively, CVaR focuses not just on the threshold where extreme losses begin, but on the full tail distribution of worse-than-expected outcomes, making it a more robust risk metric for uncertain environments.
To facilitate computational tractability, (2) can be reformulated as a linear program. Introducing an auxiliary variable $\theta$ to represent the VaR estimate, and slack variables $\pi_{\omega}$ to model excess losses in each scenario $\omega\in\Omega$, the equivalent formulation becomes:
@@
$\text{CVaR}_{\alpha}(R)=\min_{\theta,\pi}\left\{\theta+\frac{1}{1-\alpha}\sum_{\omega\in\Omega}p_{\omega}\pi_{\omega}\mid\pi_{\omega}\geq y_{\omega}-\theta,\;\pi_{\omega}\geq 0,\;\forall\omega\in\Omega\right\}$ (3)
In this expression, $\theta$ approximates the VaR at level $\alpha$, $\pi_{\omega}$ captures the excess loss in scenario $\omega$ beyond $\theta$, $p_{\omega}$ denotes the probability of scenario $\omega$ occurring, $y_{\omega}$ represents the realization of $R$ under scenario $\omega$.
This linear programming form allows CVaR minimization to be embedded efficiently into stochastic optimization models, such as our two-stage framework for intermodal logistics.
Further, CVaR admits a dual representation that highlights its interpretation as a worst-case expectation over adverse outcomes. The dual form can be expressed as:
$\text{CVaR}_{\alpha}(R)=\max_{q}\left\{\frac{1}{1-\alpha}\sum_{\omega\in\Omega}q_{\omega}y_{\omega}\mid\sum_{\omega\in\Omega}q_{\omega}=1-\alpha,\;0\leq q_{\omega}\leq p_{\omega},\;\forall\omega\in\Omega\right\}$ (4)
Alternatively, it can be interpreted through an integral form over the upper tail quantiles:
$\text{CVaR}_{\alpha}(R)=\frac{1}{1-\alpha}\int_{\alpha}^{1}\theta_{q}(R)\,dq$ (5)
where $\theta_{q}(R)$ denotes the $q$-quantile of the loss distribution.
The dual form emphasizes that CVaR aggregates weighted worst-case losses beyond the VaR threshold, providing a coherent and convex risk measure, satisfing desirable properties such as monotonicity, subadditivity, and translation equivariance (Artzner et al.,, 1999). This focus on extreme scenarios makes CVaR well-suited for applications like intermodal logistics, where demand variability and capacity fluctuations can lead to significant operational disruptions.
This dual structure enables flexible integration into optimization models, supporting resilient logistics planning under extreme uncertainties. In our work, CVaR is incorporated directly into the objective function of the two-stage stochastic optimization model, allowing for simultaneous minimization of expected transportation costs and explicit control over downside risk. This risk-aware formulation enhances the model’s ability to generate resilient, cost-effective, and environmentally sustainable operational plans. Building on this foundation, we now develop the two-stage stochastic optimization model that incorporates CVaR for intermodal freight logistics under uncertainty.
@@ 4.2.1 Objective
Min $\left(\sum_{i\in I}O_{i}y_{i}+(1-\lambda)\mathbb{E}\left(S^{\prime\prime}(a,b,\zeta(\omega),\gamma(\omega)))\right)+\lambda\text{CVaR}_{\alpha}\left(S^{\prime\prime}(a,b,\zeta(\omega),\gamma(\omega))\right)+\rho\sum_{\omega\in\Omega}p_{\omega}\eta_{\omega}\right)$ (6)
The objective function (6) integrates both the preparation (first-stage) and operational costs (second-stage) for the intermodal stochastic optimization model. This objective minimizes the mean-risk function related to the total stochastic cost. Here, ($\sum_{i\in I}O_{i}y_{i}$) represents the fixed cost of preparing containers at origin locations. The term $(1-\lambda)\mathbb{E}\left(S^{\prime\prime}(a,b,\zeta(\omega),\gamma(\omega))\right)$ reflects the expected second-stage cost, including transportation and unmet demand across all scenarios. The term $\lambda\text{CVaR}_{\alpha}\left(S^{\prime\prime}(a,b,\zeta(\omega),\gamma(\omega)))\right)$ captures tail-risk aversion, where $\lambda$ and $\alpha$ are risk parameters controlling the weight and confidence level, respectively. The final term, $\rho\sum_{\omega\in\Omega}p_{\omega}\eta_{\omega}$, penalizes emissions exceeding allowable thresholds, thereby embedding sustainability directly into the decision-making process.
@@ 4.2.2 Constraints
The constraints applied to the model effectively determine demand assignments, the number of containers on each train per hub, and limit the carbon emissions during the time period. To start with, constraints (7) ensure that the total number of containers transported from each origin $i$ does not exceed the available supply at the origin. Constraints (8) ensure that the number of containers transported to each train station does not exceed the available capacity on each train at a particular station. Further, constraints (9) ensure that container flows $x_{ijn\omega t}$ are linked to binary decision variables $z_{ijn\omega t}$, ensuring that containers can only be shipped if the corresponding binary variable is activated ($z_{ijn\omega t}=1$). The large constant $M$ acts as an upper bound, preventing any flow $x_{ijn\omega t}$ unless the route is selected.